### v1

# AUDETER: A Large-scale Dataset for Deepfake Audio Detection in Open Worlds

Abstract.
Speech generation systems can produce remarkably realistic vocalisations that are often indistinguishable from human speech, posing significant authenticity challenges. Although numerous deepfake detection methods have been developed, their effectiveness in real-world environments remains unrealiable due to the domain shift between training and test samples arising from diverse human speech and fast evolving speech synthesis systems. This is not adequately addressed by current datasets, which lack real-world application challenges with diverse and up-to-date audios in both real and deepfake categories. To fill this gap, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale, highly diverse deepfake audio dataset for comprehensive evaluation and robust development of generalised models for deepfake audio detection. It consists of over 4,500 hours of synthetic audio generated by 11 recent TTS models and 10 vocoders with a broad range of TTS/vocoder patterns, totalling 3 million audio clips, making it the largest deepfake audio dataset by scale. Through extensive experiments with AUDETER, we reveal that i) state-of-the-art (SOTA) methods trained on existing datasets struggle to generalise to novel deepfake audio samples and suffer from high false positive rates on unseen human voice, underscoring the need for a comprehensive dataset; and ii) these methods trained on AUDETER achieve highly generalised detection performance and significantly reduce detection error rate by 44.1% to 51.6%, achieving an error rate of only 4.17% on diverse cross-domain samples in the popular In-the-Wild dataset, paving the way for training generalist deepfake audio detectors. AUDETER is available on GitHub.
1. Introduction
Deepfake audio detection is the task of identifying audio generated by speech synthesis models, such as Text-To-Speech (TTS) systems and vocoders. There has been a long history of developing detection models for fake audio due to the significance of its real-world applications such as authentication in forensics, misinformation detection on social media, and voice biometric security systems.
Numerous detection methods have been developed and demonstrated to be effective when evaluated against current benchmark datasets, such as ASVSpoof (Wang et al., 2020; Yamagishi et al., 2021; Wang et al., 2024a) and In-the-Wild (Müller et al., 2022). However, the problem of open-world detection remains a major challenge, i.e., detecting deepfake audio samples generated by novel speech synthesis systems (Shen et al., 2018) that are not represented in the training samples, together with human voices with different acoustic features and artefacts. This is because most existing methods treat the detection problem as a closed-set binary classification problem and are optimised to fit the training dataset with limited audio patterns. They fail to consider potential novel patterns at inference time, which results in unsatisfactory generalisation to novel audio samples, especially when deployed in open-world detection settings. Figure 1 illustrates this limitation, showing that a state-of-the-art XLR-S-based detector (Tak et al., 2022) fails at open-world detection on two subsets of our dataset, incorrectly mapping test samples to the wrong regions due to poor generalisation to both deepfake audio samples generated by recent speech models and diverse human voices.
Current datasets have become increasingly limited in covering the diverse and up-to-date deepfake speech patterns. While existing models achieve high performance on these datasets, they do not adequately evaluate the realistic challenges for the detection in open worlds, as previously mentioned. To improve and promote evaluation for open-world detection, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale deepfake audio detection dataset that collects audio samples generated by a wide variety of speech synthesis models and multiple sources of human voices. Table 1 compares AUDETER with the existing datasets. AUDETER contains 4,682 hours of deepfake audio generated using 21 speech synthesis systems, including 10 recent Text-to-Speech (TTS) models, corresponding to 4 human voice corpora, making it much larger and more diverse than previous datasets. A key advantage is that for each real audio sample, we provide corresponding fake audio generated by all synthesis systems using matching scripts, allowing for systematic, balanced evaluation with consistent structure. By combining different AUDETER components, we can simulate various domain shifts in open-world detection.
Dataset # Audio Clips Total Hours Diverse Real # Synthetic Models # TTS Models # Identical Script ASV 2019 (Wang et al., 2020) 312K 60-70 N 19 N N ASV 2021 DF (Yamagishi et al., 2021) 612K 100-120 N 13 N N In-the-Wild (Müller et al., 2022) 31.8K 17.2 N - N N WaveFake (Frank and Schönherr, 2021) 137K 175 N 6 N Y LibriSeVoc (Sun et al., 2023) - 126.41 N 6 N Y AUDETER (ours) 3M 4,681.9 25 21 11 Y
Through extensive evaluation using AUDETER, we find that existing detection models trained on current datasets experience significant performance drops when used for open-world detection, further demonstrating the limitations in current benchmarks and training resources. Our key finding is that while generalising to all possible acoustic patterns is infeasible, speech synthesis systems share similar audio patterns. By combining a set of diverse, representative systems, we can achieve reasonably universal generalisation to most systems. In addition, training with large-scale data incorporating diverse sources of human voices will improve the domain transferability of learned deepfake patterns against domain shift in real audio. AUDETER’s consistent structure and systematic design enables understanding of the relationships between systems and human voice sources for developing more effective training data compositions.
Owing to the high diversity and large scale, AUDETER serves as a valuable resource for training open-world detectors and enables a data-centric approach to improve detection performance. All models trained on AUDETER achieved significantly improved performance compared to their pretrained versions, achieving an Equal Error Rate of 4.17% using our XLR-SLS detector. As large audio backbones become increasingly popular, the scale of our dataset can support training these data-demanding models containing millions of parameters. Our main contributions are as follows:
We identify and analyse fundamental limitations of existing deepfake audio detection methods in open-world scenarios, demonstrating through systematic evaluation that these methods are essentially closed-set binary classification approaches, failing to generalise to novel speech synthesis systems and diverse human voice characteristics not represented in training data.
We introduce AUDETER, a large-scale deepfake audio detection dataset comprising 4,682 hours of synthetic audio generated by 21 recent speech synthesis systems across 4 human voice corpora, which is substantially larger and more diverse than existing benchmarks, with systematic design enabling comprehensive evaluation of domain shifts.
We train three popular detector architectures using AUDETER and show that AUDETER effectively provides diverse audio samples as a data-centric approach for improving open-world detection, achieving superior performance compared to SOTA methods.
2. Related Work and Preliminaries
2.1. Synthetic Audio Detection
Detection models for deepfake audio have been extensively studied (Wang et al., 2020; Yamagishi et al., 2021; Frank and Schönherr, 2021; Yi et al., 2022; Müller et al., 2022; Sun et al., 2023; Wang et al., 2024a; Zhang et al., 2025; Ren et al., 2025). Earlier detection frameworks concentrated on exploring different types of features, such as short-term spectral (Xiao et al., 2015; Tian et al., 2016), long-term spectral (Alegre et al., 2013b, a; Sahidullah et al., 2015), prosodic (Pal et al., 2018; Xue et al., 2022; Wang et al., 2023), and deep features (Yu et al., 2017; Sailor et al., 2017; Cheuk et al., 2020; Fu et al., 2022; Zeghidour et al., 2018; Ravanelli and Bengio, 2018). Another direction leverages end-to-end deep neural networks for enhanced performance with a diverse range of detector architectures (Wu et al., 2020; Liu et al., 2023; Tak et al., 2021a; Jung et al., 2022; Ge et al., 2021b). More recent works employ pretrained audio models such as Wav2Vec2.0 (Baevski et al., 2020) as backbones, leveraging their substantially stronger capacity with millions of parameters and extensive pretraining knowledge in the audio domain for feature extraction, These backbones (Babu et al., 2021; Conneau et al., 2020) can then be combined in a pipeline with deep neural networks similar to the previously mentioned architectures serving as scoring networks. They have demonstrated significant performance improvements and represent the current state-of-the-art (Yang et al., 2024; Tak et al., 2022; Zhang et al., 2024). However, despite their enhanced learning capacity, they rely on supervised learning frameworks, and therefore are not readily adaptable to novel audio patterns.
2.2. Speech Synthesis Models
2.2.1. End-to-end TTS Systems
TTS systems convert text input to audio waveforms through neural networks. Earlier approaches like Tacotron (Wang et al., 2017), WaveNet (Oord et al., 2016), and FastSpeech (Ren et al., 2019) follow a two-stage workflow, in which text inputs are first converted to acoustic features, which are then processed by vocoders to generate waveforms. More sophisticated generative models are used to improve this process such as flow-based generation (Glow-TTS) (Kim et al., 2020) and end-to-end variational approaches (VITS) (Kim et al., 2021). More recent models such as YourTTS (Casanova et al., 2022) and OpenVoice (Qin et al., 2023) enable voice cloning with basic emotion and style control. However, these methods rely on phonetic alignments for generation and suffer from limited emotional realism and poor prosodic modelling capabilities such as intonation. More recent TTS systems (Du et al., 2024; Wang et al., 2025; Liao et al., 2024; Suno AI, 2023; Kim et al., 2021) leverage large language models and large-scale pretraining and significantly improved the quality and realism of speech generation.
2.2.2. Vocoders
Vocoders convert intermediate acoustic features (such as mel-spectrograms) into audio waveforms. Traditional methods (Morise et al., 2016; Degottex et al., 2011) used signal processing techniques, but neural vocoders have substantially improved audio quality. Early neural approaches (Oord et al., 2016; Mehri et al., 2016) use autoregressive generation but suffer from slow inference. Subsequent developments like (Prenger et al., 2019) and Parallel WaveGAN (Yamamoto et al., 2020) introduced parallel generation for faster synthesis. GAN-based vocoders such as MelGAN(Kumar et al., 2019), HiFi-GAN (Kong et al., 2020), and UnivNet (Jang et al., 2021) further improved efficiency and quality through adversarial training. More recent vocoders like BigVGAN (Lee et al., 2022) and Vocos (Siuzdak, 2023) have incorporated advanced architectures and training techniques to achieve state-of-the-art audio fidelity. Modern vocoders can generate audio that is nearly indistinguishable from real human speech.
2.2.3. Deepfake Audio Datasets
ASVSpoof series datasets (Wang et al., 2020; Yamagishi et al., 2021) are widely used for deepfake audio detection. ASVSpoof 2019 focuses on text-to-speech and voice conversion attacks, while ASVSpoof 2021 expanded to include more diverse spoofing methods.In-the-Wild (Müller et al., 2022) is a common choice for evaluating cross-domain detection, but its small scale and unspecified generation methods make it unsuitable for training and fine-grained evaluation. WaveFake (Frank and Schönherr, 2021) collected synthetic audio from various vocoders models, and LibriSeVoc contributed additional evaluation data by synthesising speech from LibriSpeech (Sun et al., 2023) using various vocoding techniques. However, they are relatively small-scale and primarily feature vocoders and legacy speech generation methods, with limited representation of recent end-to-end TTS systems.
2.3. Problem Statement
The objective of deepfake audio detection is to train a scoring mechanism that assigns scores to audio samples that are indicative of their authenticity. Within a normalised scale of 0 to 1, where 0 and 1 are arbitrarily set to represent fake and real, the learning scoring function should ideally satisfy , for any real audio clip and . In practice, a threshold is applied to the score as a decision boundary, classifying samples as real if and fake otherwise. The detection performance is commonly evaluated using the Equal Error Rate (EER), which is defined as the error rate at the decision threshold where the false acceptance rate (FAR) equals the false rejection rate (FRR):
where is the threshold at which the two rates are equal. A lower EER indicates better overall detection performance.
Most existing methods implement using DNNs. Open-world deepfake detection refers to detecting deepfake audio at test time when facing synthetic audio generated by novel speech synthesis systems not seen during training, as well as human voices with different styles and characteristics.
3. AUDETER Dataset
3.1. Design Motivations
The AUDETER dataset is motivated to address key limitations in current deepfake detection evaluation and bridge the gap of large-scale datasets for training generalised detection models for open-world detection.
3.1.1. Enhanced Evaluation for Open-world Detection.
Existing datasets lack comprehensive coverage of recent synthesis methods, particularly pretrained TTS models and recent vocoders that have significantly advanced quality and realism. Second, their human speech audios are limited in number and diversity. As shown in Section 4.2, current models suffer from significant performance degradation under human voice shift and novel speech synthesis systems. Therefore, we generate deepfake audio using 21 recent TTS and vocoder models to address this gap, corresponding to human speech from four corpora with different style.
3.1.2. Towards Training Generalist Detection Models.
The performance of existing models deteriorates under domain shift because they are fitted to limited training audio samples. A data-centric approach toward building generalised detection models involves training with diverse real and deepfake audio samples. This requires strong model learning capacity and large amounts of training data. AUDETER offers diverse high-quality data that can be used for large-scale training and potential self-supervised learning.
3.2. Dataset Overview
Collection Subset Partition Patterns # Audio / Patt. Total Hrs TTS In-the-Wild Bona-fide 15 19,784 311.5 Common Voice Val 15 16,372 275.0 Test 15 16,372 265.3 People Speech Val 15 18,622 493.5 Test 15 34,898 909.4 MLS Dev 15 3,807 212.7 Test 15 3,769 209.1 Vocoder In-the-Wild Bona-fide 10 19,784 207.6 Common Voice Val 10 16,372 266.7 Test 10 16,372 264.8 People Speech Val 10 18,622 331.7 Test 10 34,898 598.1 MLS Dev 10 3,807 156.7 Test 10 3,769 154.9
TTS Systems CosyVoice (2025) (Du et al., 2024), Zonos (2025) (Zyphra-Zonos, 2024), SparkTTS (2025) (Wang et al., 2025), F5-TTS (2025) (Chen et al., 2024), Fish-Speech (2024) (Liao et al., 2024), OpenVoice V2 (2023) (Qin et al., 2023), ChatTTS (2024) (2noise, 2024), XTTS v2 (2024) (Casanova et al., 2024), Bark (2023) (Suno AI, 2023), YourTTS (2022) (Casanova et al., 2022), VITS (2021) (Kim et al., 2021) Vocoders BigVGan (2022) (Lee et al., 2022), BigVSan (2024) (Shibuya et al., 2024), Vocos (Siuzdak, 2023) (2023), UnivNet (2021) (Jang et al., 2021), HiFi-GAN (2020) (Kong et al., 2020):, MelGAN (2019) (Kumar et al., 2019) , Full-band MelGAN (kan bayashi, 2019), Multi-band MelGAN (Yang et al., 2021), Parallel WaveGAN (Yamamoto et al., 2020), Style MelGAN (Mustafa et al., 2021).
We summarise the structure of AUDETER in Table 2. AUDETER consists of two collections: the TTS and the Vocoder collection, where the TTS collection contains multiple versions of deepfake audio generated using recent end-to-end TTS systems that pronounce the identical scripts as their corresponding real audio. The Vocoder collection contains vocoded audio of the real audio. Each collection is further divided into four subsets according to the source of real audio. For instance, the validation subset of the Common Voice partition consists of 16,372 real audio samples and 16 versions × 16,372 TTS-generated audio samples and 10 versions × 16,372 vocoded audio samples.
3.2.1. The Sources of Real Speech.
We include real audio samples from 4 corpora to create diverse human speech distributions. Specifically, we include all English real audio samples from the In-the-wild dataset, validation and test partitions of the Common Voice and People’s Speech Dataset, and dev and test partitions from the multilingual LibriSpeech dataset, which provides complementary characteristics. In-the-Wild offers varied recording conditions and speakers, Common Voice provides crowdsourced read speech with diverse accents, People’s Speech contains broadcast-quality professional recordings, and multilingual LibriSpeech contributes clean audiobook recordings across multiple languages, covering comprehensive human speech variability.
3.2.2. Speech Models for Audio Synthesis.
The speech models used for synthetic audio generation are summarised in Table 3. We employ 11 popular open-source TTS systems for text-to-waveform generation. For OpenVoice V2, we use 5 of its default speakers to generate 5 versions of synthetic audio for studying the effect of voice reference. We also employ 10 recent vocoders for vocoded audio. The models are selected based on their popularity and recency.
3.2.3. Visualisation of AUDETER’s Diversity.
3.3. Synthetic Audio Generation Process
Our synthetic audio is generated through two main processing pipelines: text-to-waveform synthesis via TTS systems and voice-to-voice conversion via vocoders, which constitute our two collections. For text-to-waveform generation, on the Common Voice, People’s Speech and the MLS subsets, we feed the transcripts to the TTS systems to generate corresponding synthetic speech. On the In-the-Wild dataset, we use OpenAI Whisper to generate transcripts of the bona-fide audio and then generate synthetic speech using the identical approach as above. For the vocoding collection, we apply the selected vocoders to the real audio from the subsets. It is worth noting that the synthetic patterns from the two collections differ significantly: TTS systems encode speaker references and semantic priors in the model pretraining, with many leveraging LLMs to generate semantic tokens. Vocoders encode their pretraining patterns as well as potential artefacts from the original audio.
3.4. Data Quality Assessment
To ensure the quality of our generated audio in terms of comprehensibility and naturalness indistinguishable from human speech, we throughly evaluate their intelligibility and naturalness.
3.4.1. Intelligibility Assessments
As a common quality measure for audio generation, intelligibility evaluation focuses on assessing whether the generated audio can preserve textual transcript information (Stevens et al., 2005). To make processing millions of audio files feasible, we adopt an automated evaluation approach, for which we employ a state-of-the-art Automatic Speech Recognition (ASR) model Whisper Large V3 (Radford et al., 2022) to convert generated audio clips to transcript and measure their semantic similarities with the ground truth transcript. We adopt four widely used evaluation metrics: WER Similarity (1-WER), Word Overlap, BLEU, and Exact Match, where WER focuses on transcription accuracy at the word level, Word Overlap measures lexical similarity, BLEU evaluates overall text quality, and Exact Match requires perfect transcription alignment. Figure 4 shows the WER similarity of the speech model on their generated audio for the People’s Speech Dataset. Due to space limitation, please refer to Section B for complete results for the other datasets. We found that the TTS models can produce better or at least comparable intelligibility to the original audio, and overall perform much better than the vocoders, demonstrating that the TTS models offer distinctly different and superior intelligibility patterns, showing the value of including them for comprehensive coverage.
3.4.2. Naturalness Assessment
The naturalness assessment focuses on human-like perceptual characteristics of generated audio. Mean Opinion Score (MOS) is a subjective quality metric where human listeners rate audio quality on a scale from 1 to 5 (higher is better). Due to the subjective nature of naturalness evaluation and our dataset’s large scale, we use the NISQA framework (Mittag and Möller, 2019) for automated perceptual quality prediction without requiring reference audio. The results of the average MOS score across all datasets are shown in Figure 5. Our findings here are similar to intelligibility, indicating that modern TTS systems substantially outperform vocoders in naturalness metrics, with several models approaching human-level perceptual quality.
3.5. Computational Cost
Given the scale of our dataset and varying efficiency of speech systems, we provide an estimate of our GPU usage. Dataset generation using NVIDIA A100 and H100 clusters consumed approximately 2000 GPU hours for synthetic audio and transcript generation for quality assessment.
4. Experiments
4.1. Experimental Settings
4.1.1. Baseline Methods
We employ nine popular deepfake audio detection methods with publicly available implementations, including both crafted DNN models and detection models integrating pretrained backbones with scoring heads. Among the first category, we include RawNet2 (Tak et al., 2021a), RawGAT-ST (Tak et al., 2021b), AASIST(Jung et al., 2021), PC-Dart (Ge et al., 2021a), SAMO (Ding et al., 2023), Neural Vocoder Artifacts (NVA) (Sun et al., 2023) and Purdue M2 (Ren et al., 2025). For the second category, we adapt XLS-R + RawNet + Assist (XLS+R+A) (Tak et al., 2022) and XLS+SLS (Zhang et al., 2024).
4.1.2. Evaluation Metrics
Following (Tak et al., 2021a; Müller et al., 2022; Jung et al., 2021; Tak et al., 2022; Zhang et al., 2024), we employ the popular evaluation metric, Equal Error Rate (EER), to evaluate synthetic audio detection performance. EER represents the error rate where the false positive rate intersects the false negative rate, providing a balanced measure of detection accuracy that equally weighs errors in classifying both real and synthetic audio. Lower EER values indicate superior detection performance, with 0% representing perfect detection.
4.1.3. Implementation Details
For evaluating the pretrained detection models on AUDETER, their official code implementation are used with their best pretrained checkpoints for scoring. For data-centric supervised training experiments, we use the official implementations of the selected models and use a learning rate of 1e-06 and a batch size of 128 to train all models. Given the large number of speech system patterns in our dataset, we adopt a balanced batching strategy. At each iteration, we go through all training real samples in randomised order for half of the batch size and uniformly sample all fake training samples for the other half of the batch size. Due to space limitations, please refer to the appendix for detailed implementation details.
4.1.4. Evaluation Protocol
Since AUDETER contains multiple synthetic audio versions for each real audio subset, evaluation is performed iteratively. To evaluate a subset, real audio is combined with each synthetic version in turn, repeating for all synthetic patterns. For instance, a full evaluation round for the Common Voice subset involves combining its real audio with each TTS and vocoder version, resulting in 26 runs.
4.2. Results on AUDETER as Test Data
We evaluate pretrained baseline models on AUDETER to examine their effectiveness for identifying novel deepfake patterns and diverse human speech corpora. The performance heatmap for all speech system are detailed in Figure 6. Please refer to Tables 17–24 in the appendix for the complete results on the In-the-Wild and MLS partition. We observe that the baseline methods struggle to achieve reasonable performance for most of the experiment.
Vulnerability to Novel Deepfake Patterns. The baseline methods exhibit noticeable performance degradation when directly applied to audio samples from AUDETER, which are generated by recent speech synthesis models mostly not covered by existing datasets. We found no single detection model can consistently achieve usable performance (i.e., reasonably low EER) across all datasets, indicating that generalisation achieved by training on existing datasets is insufficient for open-world detection.
We observe the following interesting phenomena. First, more recent TTS systems create greater challenges for the detectors, suggesting that the development of speech synthesis systems, particularly after the widespread adoption of pretraining, generates audio with distinct acoustic characteristics. Second, domain shift in human speech significantly affects detection performance, as performance on synthetic audio from the same TTS system varies when the corresponding real audio comes from different corpora. Third, the two baselines leveraging pretrained backbones achieved relatively better overall performance. This could be attributed to their increased parameters or pretrained knowledge enabling better characterisation of real audio. In addition, we found that baseline methods tend to perform better on deepfake audio generated by vocoders. These results demonstrate that our dataset effectively challenges existing methods and highlight the importance of evaluating against a variety of domain shifts, consistent with our motivation.
Model OV S1 OV S2 OV S2 OV S3 OV S4 RawNet2 0.632 0.235 0.572 0.211 0.225 AASIST 0.601 0.349 0.534 0.346 0.255 RawGAT-ST 0.691 0.266 0.685 0.183 0.279 PC-Dart 0.545 0.383 0.559 0.363 0.333 SAMO 0.695 0.499 0.731 0.451 0.503 NVA 0.669 0.655 0.659 0.610 0.625 Purdue M2 0.405 0.529 0.515 0.542 0.514 XLR+R+A 0.632 0.567 0.617 0.578 0.511 XLR+XLS 0.369 0.411 0.449 0.417 0.320
Effect of Speakers on Detectability. To analyse speaker effects on detection performance, we summarise baseline method EERs on synthetic audio from 5 OpenVoice versions across all datasets in Table 4. Significant variation across speakers indicates that vocal characteristics influence synthetic audio detectability beyond architectural differences.
4.3. Results on AUDETER as Training Data
AUDETER is a valuable resource for large-scale robust training. To demonstrate the effectiveness of leveraging diverse real and synthetic audio patterns as a data-centric approach for improving open-world detection performance, we perform experiments using two popular XLR-based and one DNN detection model for cross-domain evaluation. We also train the models using different combinations of AUDETER’s subsets to study the effect of training data composition. We adopt both In-the-Wild and AUDETER for open-world evaluation.
Model EER Purdue M2 (Ren et al., 2025) 79.75 PC-Dart (Ge et al., 2021a) 66.17 RawGAT-ST (Tak et al., 2021b) 52.60 AASIST (Jung et al., 2021) 43.02 RawNet2 (Tak et al., 2021a) 37.81 SAMO (Ding et al., 2023) 37.09 Wav2vec,HuBERT,Conformer & attention (Wang et al., 2023) 36.84 XLS-R & Res2Net (Yi et al., 2023) 36.62 MPE & SENet (Wang et al., 2024b) 29.62 NVA (Sun et al., 2023) 26.32 Spec & POI-Forensics (Pianese et al., 2022) 25.14 XLXS-R,WavLM,Hubert & Fusion (Yang et al., 2024) 24.27 XLR+R+A (Tak et al., 2022) 10.46 XLR-SLS (Zhang et al., 2024) 7.46 RawNet2 (Train on CV and PS subsets) 27.13 XLR+R+A (Train on CV and PS subsets) 5.05 XLR+SLS (Train on CV and PS subsets) 4.17
4.3.1. Cross Domain Generalisation from AUDETER to In-the-Wild
We choose ITW for evaluation because its synthesis methods and real audio sources are not explicitly discussed and differ from AUDETER, isolating against unintentional data leakage. We train XLS+R+A, XLR+SLS and RawNet2 using all audio from the Common Voice and People’s Speech subsets. Please note that the In-the-Wild subsets are excluded from any training combination to maintain the integrity of cross-domain evaluation.
Table 5 compares the best performance of the XLR-A-R model, XLR-SLS and RawNet 2 model trained using all data from the Common Voice and People’s Speech Dataset) with the best performance of other detection models. For comparison, we use their reported results when available, otherwise we evaluate using their official weights. The models trained using AUDETER not only achieve lower EER compared to the baseline methods, but also demonstrate significant improvement compared to the performance using their offical pretrained weights. Specifically, the ERR are reduced from 7.46 to 4.17 for XLR+SLS, 10.46 to 5.05 for SLR+R+A, and 37.81 to 27.13 for Rawnet 2, achieving respectively 44.1%, 51.6%, and 28.2% reduction in EER. This demonstrates that AUDETER provides effective diversity of deepfake and real audio combinations for improving detection model training without requiring test domain knowledge through large-scale training. The performance gain comes from diverse patterns that implicitly improve generalisation rather than data augmentation techniques. Our synthesis models, selected for popularity and recency, avoid specifically mimicking any existing benchmark.
ALL Unseen Model TTS VOC All TTS VOC All RawNet2 0.257 0.434 0.346 0.286 0.378 0.332 ASSIST 0.325 0.454 0.390 0.401 0.622 0.512 RawGAT-ST 0.266 0.475 0.371 0.346 0.447 0.397 PC-Dart 0.169 0.192 0.180 0.218 0.026 0.122 SAMO 0.492 0.506 0.499 0.471 0.475 0.473 NVA 0.492 0.551 0.522 0.450 0.408 0.429 Purdue-M2 0.438 0.564 0.501 0.537 0.715 0.626 XLR+R+A 0.365 0.192 0.279 0.346 0.239 0.292 XLR+SLS 0.299 0.148 0.224 0.275 0.164 0.220 XLR+R+A 0.035 0.018 0.026 0.028 0.015 0.021 XLR+SLS 0.016 0.013 0.014 0.010 0.026 0.018 RawNet 2 0.087 0.232 0.159 0.314 0.254 0.284
4.3.2. Cross Domain Generalisation Analysis within AUDETER
To further demonstrate that large-scale training improves generalisation, we perform cross-domain evaluation using the same three detectors trained on Common Voice and People’s Speech subsets in two settings: (1) all TTS and vocoder versions, and (2) 5 selected TTS and 5 vocoders, then evaluate on the MLS subset. The first setting examines transfer learning under human speech domain shift only, while the second examines both domain shift and novel synthesis systems. We report average performance across all system versions for the first setting and across unseen systems for the second in Table 6. We observe that all three models trained using AUDETER achieve significant EER reduction compared to all pretrained models and their pretrained versions, demonstrating that large-scale training with diverse real and synthetic audio improves generalisation under domain shift. While all three methods achieve significant overall reduction, the XLR-based models achieve near-zero performance, much better than RawNet 2, highlighting the importance of having sufficient learning capacity to incorporate open-world knowledge. Training large models requires substantial data, which is what AUDETER provides.
Collection Subset Model Best EER TTS (All) CV PS XLR+R+A 11.39 XLR-SLS 11.59 Voc (All) CV PS XLR+R+A 6.18 XLR-SLS 6.09 TTS + Voc (All) CV XLR+R+A 23.71 XLR-SLS 12.13 5 TTS + 5 Voc CV PS XLR+R+A 5.52 XLR-SLS 5.18 TTS + Vocoder (All) CV + PS XLR+R+A 5.05 XLR-SLS 4.17
4.3.3. Effect of Synthetic Pattern Diversity in Large-scale Training
To show the benefit of including diverse audio patterns in large-scale training, we vary the subsets to train multiple models and compare their performance on the in-the-wild dataset, as described in Table 7. Specifically, we train the two architecture using: (1) both TTS and Vocoder collections with only the Common Voice subset, (2) only TTS collection with Common Voice and People Speech, and (3) only Vocoder collection with Common Voice and People Speech, using the same settings. For both models, it is not surprising that the models trained using both TTS and Vocoder Collections with samples from both Common Voice and Peoples Speech partitions yield the best performance, again demonstrating the benefits of training data diversity on open-world detection performance. On the contrary, we notice that some models trained with only the Common Voice subset, even using all TTS and vocoder fake audios, produce unusable performance, further highlighting the usefulness of having diverse real audio in training.
Although models trained using the Common Voice and Peoples Speech subsets from either the TTS or the vocoder partition achieve reasonable performance, this is less than when used in combination.
4.3.4. Analysis of Training Data Scale.
To show that training with diverse data samples at greater scale can lead to better generalisation, we train the XLR+R+A model using different percentages of all data samples from the CV and PS subsets (10%, 20%, 40%, 60%, and 100%) and report the average performance of the top 3 best EER results on the In-the-Wild dataset. We observe consistent improvement as more data are used for training.
4.3.5. Single System Generalisation Test.
A key consideration in developing open-world detection models is understanding the shared characteristics between different synthesis models and how training audio from one model generalises to others. The balanced and consistent structure of AUDETER facilitates this exploration.
We train multiple XLR+R+A models using real audio from the Common Voice validation partition and synthetic audio generated by a single speech synthesis system with matching scripts. Specifically, we use deepfake samples from 5 TTS models (Fish-Speech, F5-TTS, SparkTTS, VITS, XTTS) and 5 vocoders (BigVGAN, UnivNet, MelGAN, HiFi-GAN, Vocos) and repeat the training process 10 times. Following the evaluation protocol in Sec. 4.1.4, we then evaluate the performance on different test sets. Due to space limitations, we report the average performance on the TTS and Vocoder Collection across three different test domains in Table 8. Please refer to Tables 25–36 for the complete results.
Generalisation on Matched Synthetic Audio and Identical Human Speech. The results under the CV Val column show the average performance of models trained using single systems for each collection, where scripts and real audio are identical to training, evaluating the impact of synthetic patterns alone. Generalisation ability varies significantly across speech models. For instance, SparkTTS and Fish-Speech showing stronger generalisation among TTS systems. We observe that TTS systems and vocoders struggle to generalise to each other, with performance generally better within the same collection of models. Additionally, generalisation is more difficult for vocoders. These findings demonstrate the importance of including samples from both system types and diverse training systems for robust model training.
Generalisation with Different Textual Content. We repeat the evaluation process on the Common Voice test partition, where textual information no longer matches but the same style of human voice is maintained. The results under the CV Test column shows marginally lower but comparable results compared to matched textual information, indicating that text content does not significantly affect generalisation.
Generalisation across Domains and Text. We further evaluate using the People’s Speech test partition (i.e., the PS Test column in Table 8) to explore the effect of distribution shift with different text content. Performance significantly decreases compared to previous experiments, even for the same system used in training. This suggests that detection models are sensitive to real audio distribution shift, emphasising the importance of training data diversity.
CV Val CV Test PS Test Single Sys. TTS AVG Voc Avg TTS AVG Voc Avg TTS AVG Voc Avg SparkTTS 0.096 0.496 0.100 0.496 0.482 0.541 F5-TTS 0.150 0.492 0.155 0.491 0.660 0.569 Fish-Speech 0.093 0.438 0.097 0.442 0.510 0.464 XTTS 0.172 0.496 0.178 0.496 0.669 0.531 VITS 0.108 0.485 0.111 0.485 0.455 0.533 BigVGAN 0.547 0.270 0.565 0.296 0.495 0.426 HiFi-GAN 0.238 0.285 0.249 0.290 0.691 0.441 Vocos 0.262 0.062 0.302 0.090 0.613 0.320 UnivNet 0.233 0.285 0.212 0.126 0.399 0.227 Mel GAN 0.235 0.199 0.245 0.201 0.697 0.315
5. Conclusion and Future Works
In this paper, we introduce AUDETER, a large-scale deepfake audio detection dataset for systematic benchmarking and robust large-scale training of deepfake audio detectors in open-world detection. AUDETER contains nearly 3 million synthetic audio samples generated by 21 recent speech synthesis systems that correspond to diverse selections of real human speech from 4 corpora, enabling fine-grained and balanced evaluation across human voice domains and speech systems, under various domain shifts. Through extensive experiments using the complete dataset, we demonstrate that AUDETER effectively challenges existing detection models and reveals their limitations. We also conduct large-scale training experiments using different subset combinations of AUDETER. These experiments achieve superior open-world detection performance, demonstrating AUDETER’s value as a resource for data-centric improvements and highlighting the importance of high-quality training data. We show that using high-volume diverse deepfake samples with large audio backbones effectively improves generalisation to novel audio samples.
We plan to continue developing AUDETER as an ongoing project to address the rapidly evolving nature of speech synthesis systems. We recognise that even large-scale approaches will eventually become outdated as new synthesis methods emerge. Looking forward, we plan to identify and extract representative synthesis patterns that can generalise across multiple systems to avoid continuously scaling up our dataset. Another promising direction is to explore advanced training methodologies for improved generalisation performance, such as self-supervised pretraining.
Appendix A Detailed Dataset Information
A.1. Audio Synthesis Models
We employ both recent end-to-end TTS systems and legacy vocoders for synthetic audio generation. We provide their descriptions in this section.
| Model | Link |
| CosyVoice | https://github.com/FunAudioLLM/CosyVoice |
| Zonos | https://github.com/Zyphra/Zonos |
| Spark-TTS | https://github.com/SparkAudio/Spark-TTS |
| F5-TTS | https://github.com/SWivid/F5-TTS |
| Fish-Speech | https://github.com/fishaudio/fish-speech |
| OpenVoice | https://github.com/myshell-ai/OpenVoice |
| ChatTTS | https://github.com/2noise/ChatTTS |
| Bark | https://github.com/suno-ai/bark |
| XTTS-v2 | https://github.com/coqui-ai/TTS |
| YourTTS | https://github.com/Edresson/YourTTS |
| VITS | https://github.com/jaywalnut310/vits |
A.1.1. TTS Models Selection
We include 11 popular TTS models for text to waveform deepfake audio generation. To investigate the effect of different speaker references, for OpenVoice V2, we generate 5 versions using its default speakers with 5 different English Accent accents, totalling 15 different versions. A summary of the key dataset information are reported in Table.
A.1.2. Vocoder Models Selection
We include 10 popular vocoder models for constructing the vocoder collection, and summarise their details in Table 10.
| Vocoder | Link |
| BigVGAN | https://github.com/NVIDIA/BigVGAN |
| Vocos | https://github.com/gemelo-ai/Vocos |
| BigVSAN | https://github.com/sony/bigvsan |
| UnivNet V2 | https://github.com/maum-ai/univnet |
| HiFi-GAN | https://github.com/kan-bayashi/ParallelWaveGAN |
| MelGAN | https://github.com/kan-bayashi/ParallelWaveGAN |
| MB Mel | https://github.com/kan-bayashi/ParallelWaveGAN |
| FB Mel | https://github.com/kan-bayashi/ParallelWaveGAN |
| Style Mel | https://github.com/kan-bayashi/ParallelWaveGAN |
| PW GAN | https://github.com/kan-bayashi/ParallelWaveGAN |
A.2. Real Audio
We summarise the four corpora of of real huamn voice in Table 11.
| Dataset Name | Link |
| In-the-Wild | https://deepfake-total.com/in_the_wild |
| Common Voice 13. 0 | Hugging Face Datasets: mozilla-foundation/common_voice_13_0 (validation and test partition) |
| The People’s Speech | Hugging Face Datasets: MLCommons/peoples_speech (validation and test partition from the clean subset) |
| Multilingual LibriSpeech (MLS) | Hugging Face Datasets: parler-tts/mls_eng (English version of the Multilingual LibriSpeech (MLS) dataset) |
A.3. Detailed Dataset Statistics
We present the detailed hour counts for each partition of the TTS collection in Table 12 and the vocoder collection in Table 13.
| TTS Collection | ||||||||||
| In-the-Wild | Common Voice | People Speech | MLS | |||||||
| Model | Bona-fide | Val | Test | Total | Val | Test | Total | Dev | Test | Total |
| CosyVoice | 16.25 | 14.76 | 14.00 | 28.76 | 27.15 | 50.04 | 77.19 | 12.48 | 12.45 | 24.93 |
| Zonos | 18.01 | 16.20 | 15.48 | 31.68 | 30.09 | 55.50 | 85.59 | 12.94 | 12.96 | 25.90 |
| Sparktts | 18.89 | 18.24 | 17.44 | 35.68 | 30.87 | 56.88 | 87.75 | 13.95 | 13.91 | 27.86 |
| F5-TTS | 27.91 | 25.58 | 24.49 | 50.07 | 48.83 | 89.83 | 138.66 | 21.75 | 21.68 | 43.43 |
| Fish-Speech | 27.03 | 18.07 | 18.70 | 36.77 | 40.70 | 74.27 | 114.97 | 33.75 | 30.55 | 64.30 |
| OV2 S1 | 20.74 | 18.92 | 18.20 | 37.12 | 31.16 | 57.62 | 88.78 | 12.79 | 12.77 | 25.56 |
| OV2 S2 | 17.28 | 15.90 | 15.23 | 31.13 | 27.37 | 50.56 | 77.93 | 11.55 | 11.54 | 23.09 |
| OV2 S3 | 19.53 | 17.93 | 17.24 | 35.17 | 30.04 | 55.52 | 85.56 | 12.40 | 12.39 | 24.79 |
| OV2 S4 | 19.98 | 18.02 | 17.36 | 35.38 | 29.98 | 55.39 | 85.37 | 12.07 | 12.06 | 24.13 |
| OV2 S5 | 16.02 | 14.74 | 14.09 | 28.83 | 25.61 | 47.29 | 72.90 | 11.04 | 11.04 | 22.08 |
| ChatTTS | 24.95 | 21.73 | 20.98 | 42.71 | 44.33 | 81.13 | 125.46 | 20.49 | 20.28 | 40.77 |
| Bark | 29.05 | 24.79 | 24.42 | 49.21 | 40.18 | 74.26 | 114.44 | 13.96 | 13.79 | 27.75 |
| XTTS | 18.54 | 16.37 | 15.71 | 32.08 | 30.88 | 57.05 | 87.93 | 12.78 | 12.77 | 25.55 |
| YourTTS | 20.14 | 18.56 | 17.76 | 36.32 | 30.88 | 57.05 | 87.93 | 12.78 | 12.84 | 25.62 |
| VITS | 17.16 | 15.21 | 14.17 | 29.38 | 25.42 | 47.01 | 72.43 | 10.49 | 10.52 | 21.01 |
| Total | 311.48 | 275.02 | 265.27 | 540.29 | 493.49 | 909.40 | 1402.89 | 225.22 | 221.55 | 446.77 |
| Vocoder Collection | ||||||||||
| In-the-Wild | Common Voice | People Speech | MLS | |||||||
| Model | Bona-fide | Val | Test | Total | Val | Test | Total | Val | Test | Total |
| BigVGAN | 20.73 | 27.24 | 27.04 | 54.28 | 33.14 | 59.75 | 92.89 | 15.75 | 15.54 | 31.29 |
| Vocos | 20.73 | 27.24 | 27.04 | 54.28 | 33.14 | 59.75 | 92.89 | 15.75 | 15.54 | 31.29 |
| BigVSAN | 20.73 | 27.16 | 26.96 | 54.12 | 33.14 | 59.75 | 92.89 | 15.76 | 15.55 | 31.31 |
| UnivNet | 20.73 | 25.88 | 25.71 | 51.59 | 33.14 | 59.75 | 92.89 | 15.42 | 15.42 | 30.84 |
| HiFi-GAN | 20.79 | 25.88 | 25.71 | 51.59 | 33.19 | 59.86 | 93.05 | 15.59 | 15.42 | 31.01 |
| MelGAN | 20.77 | 25.88 | 25.71 | 51.59 | 33.19 | 59.85 | 93.04 | 15.59 | 15.42 | 31.01 |
| MB Mel | 20.77 | 27.16 | 26.96 | 54.12 | 33.19 | 59.85 | 93.04 | 15.76 | 15.55 | 31.31 |
| FB Mel | 20.77 | 25.88 | 25.71 | 51.59 | 33.19 | 59.85 | 93.04 | 15.59 | 15.42 | 31.01 |
| Style Mel | 20.79 | 27.24 | 27.04 | 54.28 | 33.19 | 59.86 | 93.05 | 15.75 | 15.54 | 31.29 |
| PW GAN | 20.77 | 27.11 | 26.91 | 54.02 | 33.19 | 59.85 | 93.04 | 15.75 | 15.54 | 31.29 |
| Total | 207.58 | 266.67 | 264.79 | 531.46 | 331.7 | 598.12 | 929.82 | 156.71 | 154.94 | 311.65 |
Appendix B Complete Datasets Quality Assurance Results
We provide the detailed results for intelligibility assessment and naturalness.
| Metricc | Original | CosyVoice | Zonos | Spark-TTS | F5-TTS | Fish-Speech | O.V S1 | O.V S2 | O.V S3 | O.V S4 | O.V S5 | ChatTTS | Bark | XTTS-v2 | YourTTS | VITS |
| WER | 0.873 | 0.935 | 0.869 | 0.962 | 0.920 | 0.924 | 0.950 | 0.898 | 0.950 | 0.953 | 0.955 | 0.887 | 0.850 | 0.946 | 0.894 | 0.908 |
| Word Overlap | 0.846 | 0.905 | 0.826 | 0.943 | 0.924 | 0.888 | 0.924 | 0.858 | 0.924 | 0.929 | 0.932 | 0.853 | 0.842 | 0.921 | 0.853 | 0.874 |
| BLEU | 0.816 | 0.887 | 0.799 | 0.931 | 0.912 | 0.866 | 0.910 | 0.833 | 0.910 | 0.915 | 0.917 | 0.833 | 0.823 | 0.910 | 0.834 | 0.850 |
| Exact Match | 0.545 | 0.659 | 0.469 | 0.780 | 0.705 | 0.613 | 0.712 | 0.542 | 0.718 | 0.734 | 0.722 | 0.519 | 0.509 | 0.682 | 0.510 | 0.592 |
| Metricc | Original | CosyVoice | Zonos | Spark-TTS | F5-TTS | Fish-Speech | O.V S1 | O.V S2 | O.V S3 | O.V S4 | O.V S5 | ChatTTS | Bark | XTTS-v2 | YourTTS | VITS |
| WER | 0.707 | 0.912 | 0.853 | 0.943 | 0.859 | 0.921 | 0.931 | 0.906 | 0.910 | 0.944 | 0.944 | 0.863 | 0.805 | 0.924 | 0.888 | 0.907 |
| Word Overlap | 0.765 | 0.891 | 0.827 | 0.931 | 0.912 | 0.895 | 0.906 | 0.880 | 0.886 | 0.924 | 0.925 | 0.848 | 0.832 | 0.913 | 0.864 | 0.879 |
| BLEU | 0.724 | 0.873 | 0.798 | 0.917 | 0.891 | 0.871 | 0.887 | 0.857 | 0.864 | 0.909 | 0.909 | 0.817 | 0.802 | 0.896 | 0.844 | 0.850 |
| Exact Match | 0.239 | 0.562 | 0.402 | 0.688 | 0.584 | 0.540 | 0.583 | 0.525 | 0.256 | 0.650 | 0.650 | 0.411 | 0.405 | 0.603 | 0.501 | 0.474 |
B.1. Detailed Naturalness Scores
We provide the details results for the NISQA naturalness assessmenton all our deepfake audio samples and present the results in Table 16.
| itw | cv | ps | mls | ||||
| bona-fide | val | test | val | test | dev | test | |
| CosyVoice | 4.75 | 4.817 | 4.81 | 4.781 | 4.771 | 4.806 | 4.805 |
| Zonos | 4.736 | 4.856 | 4.844 | 4.805 | 4.798 | 4.848 | 4.859 |
| Spark-TTS | 4.752 | 4.797 | 4.79 | 4.739 | 4.736 | 4.747 | 4.742 |
| F5-TTS | 3.908 | 4.163 | 4.128 | 3.981 | 3.997 | 4.053 | 4.061 |
| Fish-Speech | 4.353 | 4.379 | 4.371 | 4.423 | 4.41 | 4.491 | 4.493 |
| O.V s1 | 4.466 | 4.506 | 4.51 | 4.548 | 4.534 | 4.619 | 4.639 |
| O.V s2 | 4.566 | 4.622 | 4.619 | 4.53 | 4.524 | 4.537 | 4.534 |
| O.V s3 | 4.977 | 5.007 | 5.007 | 5.005 | 5.002 | 5.011 | 5.008 |
| O.V s4 | 4.825 | 4.894 | 4.89 | 4.851 | 4.85 | 4.849 | 4.857 |
| O.V s5 | 4.521 | 4.636 | 4.633 | 4.589 | 4.577 | 4.611 | 4.6 |
| ChatTTS | 4.264 | 4.466 | 4.451 | 4.359 | 4.371 | 4.378 | 4.373 |
| Bark | 3.352 | 3.562 | 3.521 | 3.434 | 3.429 | 3.69 | 3.652 |
| XTTS | 4.302 | 4.441 | 4.414 | 4.413 | 4.398 | 4.604 | 4.601 |
| YourTTS | 4.29 | 4.487 | 4.469 | 4.474 | 4.457 | 4.603 | 4.601 |
| VITS | 4.677 | 4.778 | 0.468 | 4.772 | 4.767 | 4.824 | 4.824 |
| BigVGAN | 3.287 | 3.105 | 2.945 | 2.081 | 2.291 | 3.694 | 3.805 |
| Vocos | 3.27 | 3.079 | 2.92 | 2.088 | 2.273 | 3.633 | 3.738 |
| BigVSAN | 3.267 | 3.118 | 2.955 | 2.19 | 2.406 | 3.731 | 3.819 |
| UnivNet | 3.359 | 3.437 | 3.27 | 2.34 | 2.443 | 3.849 | 3.902 |
| HiFi-GAN | 3.014 | 2.995 | 2.865 | 1.952 | 2.074 | 3.45 | 3.572 |
| MelGAN | 2.987 | 2.941 | 2.801 | 2.226 | 2.3 | 3.221 | 3.252 |
| MB Mel | 2.874 | 3.043 | 2.92 | 2.052 | 2.116 | 3.393 | 3.383 |
| FB Mel | 3.137 | 3.267 | 3.134 | 2.22 | 2.299 | 3.634 | 3.626 |
| Style Mel | 2.826 | 2.913 | 2.801 | 1.922 | 2.036 | 3.35 | 3.421 |
| PW GAN | 2.977 | 3.193 | 3.062 | 2.082 | 2.226 | 3.547 | 3.541 |
Appendix C Complete Experimental Results
C.1. Detailed Zero-shot Performance
We present selected results of open-world detection performance of the baseline methods on the Common Voice and Peoples Speech subsets for both the TTS and Vocoder collections due to space limitations. Here, we provide the detailed results for all subsets on both collections.
model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg XLS+R+A 0.436 0.297 0.565 0.676 0.163 0.566 0.367 0.491 0.411 0.293 0.180 0.180 0.200 0.135 0.469 0.362 XLR+SLS 0.464 0.279 0.545 0.672 0.150 0.640 0.383 0.557 0.423 0.253 0.133 0.171 0.118 0.101 0.475 0.358 RawNet2 0.220 0.470 0.252 0.716 0.447 0.820 0.314 0.867 0.260 0.395 0.550 0.522 0.287 0.127 0.375 0.441 ASSIST 0.380 0.520 0.382 0.885 0.434 0.774 0.427 0.776 0.465 0.366 0.615 0.487 0.484 0.261 0.501 0.517 RawGAT-ST 0.152 0.508 0.293 0.879 0.451 0.755 0.388 0.754 0.339 0.306 0.574 0.489 0.218 0.246 0.489 0.456 PC-Dart 0.845 0.696 0.477 0.829 0.709 0.804 0.788 0.718 0.702 0.743 0.831 0.730 0.581 0.674 0.610 0.716 SAMO 0.690 0.664 0.532 0.523 0.541 0.671 0.655 0.680 0.708 0.697 0.575 0.513 0.742 0.348 0.800 0.623 NVA 0.590 0.601 0.552 0.568 0.519 0.592 0.598 0.569 0.593 0.607 0.344 0.334 0.471 0.224 0.612 0.518 Purdue-M2 0.526 0.353 0.436 0.564 0.435 0.178 0.492 0.435 0.519 0.429 0.507 0.828 0.342 0.697 0.420 0.477
Model BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg XLS+R+A 0.3820 0.2341 0.3150 0.1910 0.1398 0.0541 0.0656 0.0703 0.0675 0.066 0.159 XLR+SLS 0.3827 0.2303 0.3147 0.1661 0.1196 0.0347 0.0476 0.0516 0.0520 0.0491 0.145 RawNet2 0.4884 0.4861 0.5022 0.4565 0.4299 0.5224 0.4819 0.4124 0.3846 0.4769 0.464 ASSIST 0.5089 0.5064 0.5362 0.5667 0.4773 0.4564 0.4515 0.4062 0.3811 0.5013 0.479 RawGAT-ST 0.5083 0.4982 0.5256 0.5653 0.4821 0.5244 0.4990 0.4596 0.3909 0.5417 0.500 PC-Dart 0.7381 0.7805 0.8424 0.7005 0.6718 0.6968 0.6407 0.7457 0.6129 0.6756 0.711 SAMO 0.5081 0.5012 0.5302 0.6079 0.5262 0.5193 0.5350 0.4728 0.4531 0.5497 0.520 NVA 0.5046 0.4929 0.5002 0.5254 0.5230 0.5004 0.4639 0.4710 0.5141 0.5292 0.502 Purdue-M2 0.5101 0.5925 0.5303 0.5018 0.5851 0.7995 0.8013 0.7372 0.7481 0.7258 0.653
model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg XLS+R+A 0.343 0.207 0.786 0.706 0.078 0.499 0.225 0.402 0.244 0.142 0.083 0.096 0.084 0.041 0.381 0.288 XLR+SLS 0.418 0.261 0.642 0.600 0.134 0.524 0.337 0.455 0.336 0.202 0.101 0.135 0.083 0.061 0.396 0.312 RawNet2 0.188 0.417 0.600 0.705 0.383 0.756 0.271 0.808 0.172 0.316 0.491 0.424 0.220 0.097 0.288 0.409 ASSIST 0.451 0.531 0.661 0.928 0.484 0.721 0.479 0.738 0.474 0.429 0.633 0.462 0.503 0.358 0.490 0.556 RawGAT-ST 0.157 0.509 0.634 0.898 0.456 0.684 0.394 0.728 0.279 0.294 0.548 0.424 0.198 0.232 0.431 0.458 PC-Dart 0.634 0.432 0.306 0.600 0.451 0.546 0.501 0.597 0.409 0.424 0.640 0.481 0.442 0.648 0.451 0.504 SAMO 0.777 0.728 0.629 0.579 0.624 0.771 0.739 0.770 0.763 0.781 0.661 0.514 0.819 0.410 0.867 0.695 NVA 0.708 0.715 0.702 0.690 0.670 0.714 0.719 0.710 0.718 0.721 0.450 0.433 0.635 0.377 0.713 0.645 Purdue-M2 0.316 0.175 0.163 0.289 0.228 0.064 0.257 0.188 0.254 0.229 0.274 0.687 0.170 0.481 0.199 0.265
Model BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg XLS+R+A 0.4100 0.2722 0.3410 0.1935 0.1077 0.0993 0.1224 0.1132 0.0783 0.1320 0.187 XLR+SLS 0.4891 0.5169 0.3410 0.1752 0.4915 0.4643 0.5310 0.5081 0.3636 0.4389 0.432 RawNet2 0.5053 0.4956 0.5136 0.5398 0.4989 0.4415 0.4896 0.5139 0.5241 0.4875 0.501 ASSIST 0.4929 0.5129 0.4842 0.393 0.465 0.4862 0.4965 0.5464 0.5222 0.4365 0.484 RawGAT-ST 0.4954 0.5314 0.4871 0.3842 0.4647 0.4481 0.4655 0.5210 0.4827 0.4083 0.469 PC-Dart 0.5002 0.4940 0.5383 0.4697 0.5235 0.4429 0.4263 0.4631 0.4150 0.4292 0.470 SAMO 0.5072 0.4901 0.5124 0.6392 0.5836 0.5808 0.611 0.5127 0.5150 0.6512 0.560 NVA 0.4867 0.4675 0.4862 0.5747 0.6152 0.5873 0.5561 0.5972 0.6593 0.6149 0.565 Purdue-M2 0.4847 0.5481 0.3067 0.3296 0.5649 0.7629 0.7532 0.7278 0.7001 0.7107 0.589
model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg XLS+R+A 0.3253 0.554 0.833 0.871 0.339 0.748 0.607 0.723 0.651 0.469 0.380 0.296 0.349 0.255 0.753 0.544 XLR+SLS 0.7288 0.536 0.769 0.779 0.377 0.758 0.611 0.732 0.616 0.451 0.363 0.302 0.246 0.214 0.694 0.545 RawNet2 0.2093 0.440 0.205 0.686 0.412 0.775 0.264 0.761 0.195 0.326 0.521 0.461 0.232 0.103 0.284 0.392 ASSIST 0.5022 0.626 0.448 0.945 0.538 0.840 0.497 0.771 0.504 0.435 0.709 0.528 0.573 0.298 0.524 0.582 RawGAT-ST Chicken 0.591 0.303 0.936 0.568 0.817 0.413 0.703 0.323 0.357 0.692 0.557 0.224 0.270 0.515 0.519 PC-Dart 0.0483 0.053 0.018 0.037 0.069 0.050 0.026 0.051 0.017 0.022 0.237 0.091 0.031 0.085 0.028 0.058 SAMO 0.7851 0.703 0.599 0.567 0.610 0.741 0.704 0.755 0.707 0.737 0.657 0.566 0.789 0.382 0.833 0.676 NVA 0.7231 0.694 0.660 0.660 0.597 0.695 0.699 0.685 0.694 0.703 0.401 0.415 0.616 0.378 0.705 0.622 Purdue-M2 0.4867 0.274 0.440 0.545 0.363 0.108 0.477 0.409 0.497 0.398 0.478 0.812 0.305 0.699 0.381 0.445
Model BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg XLS+R+A 0.428 0.357 0.396 0.250 0.223 0.171 0.208 0.197 0.202 0.191 0.262 XLR+SLS 0.417 0.318 0.380 0.180 0.152 0.098 0.127 0.102 0.104 0.108 0.199 RawNet2 0.464 0.461 0.468 0.352 0.351 0.430 0.408 0.341 0.276 0.396 0.395 ASSIST 0.497 0.501 0.494 0.606 0.643 0.649 0.637 0.689 0.709 0.582 0.600 RawGAT-ST 0.506 0.492 0.512 0.473 0.428 0.443 0.433 0.425 0.376 0.492 0.458 PC-Dart 0.028 0.032 0.036 0.024 0.026 0.024 0.023 0.028 0.021 0.022 0.026 SAMO 0.505 0.493 0.511 0.462 0.445 0.500 0.489 0.418 0.453 0.502 0.478 NVA 0.493 0.492 0.487 0.407 0.408 0.382 0.354 0.364 0.401 0.433 0.422 Purdue-M2 0.500 0.538 0.504 0.539 0.650 0.812 0.765 0.745 0.792 0.769 0.661
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg XLS+R+A 0.542 0.382 0.600 0.630 0.268 0.472 0.344 0.467 0.348 0.247 0.220 0.166 0.198 0.138 0.459 0.365 XLR+SLS 0.542 0.263 0.514 0.598 0.208 0.474 0.284 0.441 0.280 0.171 0.126 0.103 0.080 0.044 0.363 0.299 RawNet2 0.069 0.244 0.043 0.627 0.420 0.708 0.079 0.588 0.017 0.151 0.449 0.320 0.056 0.008 0.076 0.257 ASSIST 0.262 0.324 0.163 0.934 0.467 0.517 0.198 0.385 0.149 0.138 0.534 0.276 0.303 0.078 0.156 0.325 RawGAT-ST 0.087 0.307 0.094 0.875 0.597 0.503 0.086 0.336 0.019 0.112 0.467 0.315 0.041 0.024 0.133 0.266 PC-Dart 0.194 0.125 0.081 0.152 0.237 0.140 0.125 0.126 0.105 0.112 0.390 0.176 0.154 0.270 0.149 0.169 SAMO 0.682 0.527 0.349 0.206 0.465 0.571 0.473 0.674 0.424 0.630 0.477 0.346 0.720 0.120 0.724 0.492 NVA 0.604 0.595 0.572 0.566 0.504 0.611 0.611 0.610 0.610 0.611 0.289 0.365 0.554 0.281 0.611 0.533 Purdue-M2 0.440 0.322 0.412 0.484 0.424 0.114 0.464 0.420 0.445 0.349 0.426 0.829 0.343 0.723 0.381 0.438
Model BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg XLS+R+A 0.389 0.243 0.344 0.198 0.156 0.140 0.127 0.116 0.088 0.126 0.192 XLR+SLS 0.382 0.208 0.338 0.159 0.122 0.043 0.050 0.061 0.052 0.061 0.148 RawNet2 0.475 0.424 0.488 0.436 0.438 0.461 0.427 0.369 0.373 0.455 0.434 ASSIST 0.503 0.474 0.506 0.531 0.489 0.404 0.404 0.388 0.405 0.432 0.454 RawGAT-ST 0.517 0.485 0.517 0.498 0.494 0.493 0.446 0.412 0.418 0.473 0.475 PC-Dart 0.217 0.238 0.253 0.195 0.202 0.163 0.157 0.177 0.164 0.151 0.192 SAMO 0.489 0.495 0.493 0.501 0.503 0.510 0.538 0.500 0.477 0.553 0.506 NVA 0.506 0.512 0.521 0.542 0.574 0.562 0.548 0.570 0.590 0.584 0.551 Purdue-M2 0.462 0.507 0.476 0.470 0.477 0.697 0.681 0.627 0.631 0.612 0.564
C.2. Single System Generalisation Result
In this section, we present the full single system generalisation results.
C.2.1. Same textual information, same domain.
This section provides complete results for Section 4.3.5.
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg Spark-TTS 0.006 0.000 0.000 0.024 0.116 0.314 0.027 0.400 0.000 0.004 0.524 0.019 0.000 0.000 0.001 0.096 F5_tts 0.219 0.003 0.082 0.000 0.209 0.288 0.116 0.444 0.032 0.290 0.442 0.028 0.003 0.008 0.079 0.150 fish_speech 0.005 0.001 0.171 0.427 0.000 0.252 0.001 0.190 0.000 0.000 0.351 0.002 0.000 0.000 0.000 0.093 XTTS 0.112 0.007 0.242 0.517 0.229 0.292 0.035 0.318 0.068 0.097 0.581 0.055 0.000 0.008 0.024 0.172 VITS 0.004 0.000 0.071 0.349 0.090 0.244 0.002 0.279 0.000 0.001 0.574 0.012 0.000 0.000 0.000 0.108
| Train Patt. | BigVGAN | Vocos | BigVSAN | UnivNet | HiFi-GAN | MelGAN | MB Mel | FB Mel | Style Mel | PW GAN | Avg |
| Spark-TTS | 0.501 | 0.502 | 0.499 | 0.497 | 0.479 | 0.498 | 0.502 | 0.497 | 0.477 | 0.509 | 0.496 |
| F5-TTS | 0.498 | 0.496 | 0.492 | 0.499 | 0.470 | 0.498 | 0.496 | 0.495 | 0.463 | 0.513 | 0.492 |
| Fish-speech | 0.493 | 0.487 | 0.489 | 0.462 | 0.359 | 0.410 | 0.435 | 0.431 | 0.369 | 0.449 | 0.438 |
| XTTS | 0.499 | 0.499 | 0.498 | 0.506 | 0.510 | 0.487 | 0.491 | 0.488 | 0.484 | 0.496 | 0.496 |
| VITS | 0.500 | 0.500 | 0.498 | 0.501 | 0.446 | 0.483 | 0.486 | 0.483 | 0.454 | 0.500 | 0.485 |
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg Mel GAN 0.678 0.606 0.675 0.435 0.465 0.411 0.588 0.541 0.666 0.629 0.148 0.559 0.640 0.647 0.522 0.547 HiFi-GAN 0.261 0.001 0.232 0.582 0.205 0.428 0.316 0.446 0.221 0.324 0.527 0.003 0.001 0.000 0.025 0.238 Vocos 0.277 0.274 0.386 0.104 0.106 0.305 0.358 0.494 0.478 0.393 0.050 0.106 0.209 0.062 0.321 0.262 UnivNet 0.186 0.001 0.232 0.582 0.205 0.428 0.316 0.446 0.221 0.324 0.527 0.003 0.001 0.000 0.025 0.233 Mel GAN 0.221 0.134 0.282 0.441 0.200 0.359 0.188 0.433 0.170 0.281 0.302 0.150 0.095 0.107 0.158 0.235
Train Patt. BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg Spark-TTS 0.250 0.241 0.183 0.130 0.400 0.262 0.276 0.287 0.341 0.328 0.270 F5-TTS 0.490 0.476 0.479 0.433 0.000 0.277 0.232 0.195 0.019 0.245 0.285 Fish-speech 0.273 0.016 0.119 0.024 0.063 0.018 0.021 0.024 0.031 0.031 0.062 XTTS 0.490 0.476 0.479 0.433 0.000 0.277 0.232 0.195 0.019 0.245 0.285 VITS 0.489 0.467 0.477 0.351 0.163 0.000 0.006 0.015 0.012 0.012 0.199
C.2.2. Same domain different textual information.
This section provides complete results for Section 4.3.5.
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg Spark-TTS 0.007 0.001 0.001 0.030 0.126 0.329 0.032 0.407 0.001 0.005 0.528 0.021 0.001 0.001 0.005 0.100 F5_tts 0.229 0.004 0.091 0.000 0.219 0.300 0.128 0.453 0.035 0.296 0.446 0.029 0.004 0.008 0.080 0.155 Fish-Speech 0.006 0.001 0.176 0.437 0.001 0.268 0.001 0.195 0.000 0.001 0.363 0.003 0.000 0.000 0.006 0.097 XTTS 0.118 0.008 0.246 0.519 0.240 0.304 0.040 0.327 0.077 0.105 0.586 0.061 0.000 0.007 0.027 0.178 VITS 0.005 0.001 0.076 0.351 0.101 0.255 0.003 0.282 0.001 0.002 0.572 0.015 0.000 0.000 0.001 0.111
| Model | BigVGAN | Vocos | BigVSAN | UnivNet | HiFi-GAN | MelGAN | MB Mel | FB Mel | Style Mel | PW GAN | Avg |
| Spark-TTS | 0.500 | 0.503 | 0.499 | 0.497 | 0.483 | 0.496 | 0.503 | 0.495 | 0.473 | 0.511 | 0.496 |
| F5-TTS | 0.496 | 0.495 | 0.491 | 0.498 | 0.474 | 0.500 | 0.496 | 0.495 | 0.457 | 0.513 | 0.491 |
| Fish-Speech | 0.491 | 0.489 | 0.489 | 0.466 | 0.366 | 0.419 | 0.440 | 0.440 | 0.365 | 0.455 | 0.442 |
| XTTS | 0.499 | 0.501 | 0.498 | 0.507 | 0.508 | 0.489 | 0.491 | 0.486 | 0.481 | 0.500 | 0.496 |
| VITS | 0.500 | 0.502 | 0.500 | 0.500 | 0.449 | 0.484 | 0.489 | 0.484 | 0.447 | 0.501 | 0.485 |
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg BigVGAN 0.695 0.625 0.687 0.450 0.489 0.433 0.613 0.569 0.677 0.635 0.157 0.579 0.662 0.669 0.531 0.565 HiFi-GAN 0.277 0.002 0.244 0.587 0.219 0.453 0.340 0.464 0.234 0.345 0.544 0.004 0.001 0.000 0.029 0.249 Vocos 0.320 0.325 0.428 0.140 0.148 0.336 0.402 0.522 0.514 0.433 0.084 0.146 0.261 0.099 0.370 0.302 UnivNet 0.215 0.144 0.421 0.127 0.093 0.437 0.237 0.281 0.170 0.299 0.044 0.182 0.085 0.105 0.336 0.212 Mel GAN 0.227 0.147 0.287 0.447 0.213 0.371 0.202 0.439 0.181 0.284 0.316 0.163 0.112 0.120 0.164 0.245
| Train Patt. | BigVGAN | Vocos | BigVSAN | UnivNet | HiFi-GAN | MelGAN | MB Mel | FB Mel | Style Mel | PW GAN | Avg |
| BigVGAN | 0.281 | 0.274 | 0.211 | 0.148 | 0.419 | 0.292 | 0.304 | 0.314 | 0.370 | 0.350 | 0.296 |
| HiFi-GAN | 0.491 | 0.479 | 0.483 | 0.438 | 0.001 | 0.292 | 0.239 | 0.203 | 0.019 | 0.252 | 0.290 |
| Vocos | 0.287 | 0.049 | 0.153 | 0.050 | 0.099 | 0.041 | 0.045 | 0.049 | 0.063 | 0.065 | 0.090 |
| UnivNet | 0.435 | 0.324 | 0.260 | 0.012 | 0.101 | 0.012 | 0.032 | 0.033 | 0.025 | 0.028 | 0.126 |
| Mel GAN | 0.486 | 0.466 | 0.474 | 0.345 | 0.176 | 0.002 | 0.010 | 0.020 | 0.016 | 0.016 | 0.201 |
C.2.3. Different domain and different textual information.
This section provides complete results for Section 4.3.5.
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg Spark-TTS 0.301 0.307 0.320 0.542 0.473 0.998 0.379 0.998 0.246 0.336 0.979 0.507 0.332 0.291 0.217 0.482 F5_tts 0.844 0.341 0.715 0.004 0.611 0.987 0.918 0.991 0.757 0.981 0.943 0.582 0.208 0.278 0.740 0.660 fish_speech 0.400 0.359 0.459 0.948 0.372 0.980 0.334 0.878 0.277 0.291 0.935 0.463 0.370 0.358 0.232 0.510 XTTS 0.684 0.286 0.656 0.985 0.599 0.992 0.895 0.992 0.670 0.980 0.993 0.533 0.053 0.205 0.519 0.669 VITS 0.301 0.270 0.375 0.938 0.464 0.998 0.188 0.989 0.129 0.162 0.991 0.503 0.202 0.246 0.067 0.455
Model BigVGAN Vocos BigVSAN UnivNet HiFi-GAN MelGAN MB Mel FB Mel Style Mel PW GAN Avg sparktts 0.527 0.532 0.522 0.516 0.533 0.573 0.557 0.570 0.531 0.554 0.541 f5_tts 0.569 0.568 0.561 0.537 0.559 0.590 0.572 0.574 0.565 0.598 0.569 fish_speech 0.474 0.469 0.470 0.458 0.477 0.440 0.473 0.493 0.440 0.447 0.464 xtts 0.547 0.551 0.534 0.494 0.519 0.537 0.540 0.551 0.509 0.524 0.531 vits 0.5263 0.5264 0.5179 0.5064 0.5388 0.5454 0.5489 0.5591 0.5188 0.5422 0.533
Model CosyVoice Zonos Spark-TTS F5-TTS Fish-Speech O.V s1 O.V s2 O.V s3 O.V s4 O.V s5 ChatTTS Bark XTTS YourTTS VITS Avg BigVGAN 0.608 0.537 0.673 0.351 0.398 0.325 0.537 0.523 0.647 0.622 0.119 0.458 0.557 0.563 0.508 0.495 F5_tts 0.644 0.483 0.652 0.982 0.528 0.999 0.900 0.999 0.550 0.917 0.989 0.522 0.432 0.381 0.387 0.691 fish_speech 0.689 0.741 0.858 0.337 0.369 0.585 0.857 0.813 0.874 0.833 0.239 0.380 0.531 0.256 0.832 0.613 XTTS 0.458 0.288 0.704 0.283 0.210 0.711 0.481 0.479 0.353 0.551 0.108 0.358 0.175 0.226 0.606 0.399 VITS 0.726 0.669 0.751 0.855 0.579 0.826 0.695 0.844 0.675 0.775 0.639 0.586 0.573 0.586 0.671 0.697
| Model | BigVGAN | Vocos | BigVSAN | UnivNet | HiFi-GAN | MelGAN | MB Mel | FB Mel | Style Mel | PW GAN | Avg |
| BigVGAN | 0.440 | 0.401 | 0.377 | 0.343 | 0.409 | 0.426 | 0.382 | 0.383 | 0.395 | 0.411 | 0.426 |
| HiFi-GAN | 0.486 | 0.468 | 0.471 | 0.434 | 0.444 | 0.365 | 0.407 | 0.436 | 0.383 | 0.384 | 0.441 |
| Vocos | 0.398 | 0.217 | 0.301 | 0.240 | 0.229 | 0.216 | 0.205 | 0.215 | 0.207 | 0.227 | 0.320 |
| UnivNet | 0.427 | 0.346 | 0.272 | 0.077 | 0.146 | 0.070 | 0.093 | 0.104 | 0.105 | 0.087 | 0.227 |
| Mel GAN | 0.485 | 0.423 | 0.456 | 0.290 | 0.149 | 0.037 | 0.065 | 0.083 | 0.084 | 0.065 | 0.315 |